Mr. Javed approached Islamic Bank to purchase a house built at DHA Phase VIII on 1,000 sq yards. The total value of the house is Rs. 25 Million. He only has 04 Million and wants Islamic Bank to invest the rest, which Islamic Bank agrees. 

 Mr. Javed agrees to purchase the Islamic Banks’s share in the house in 05 years' time on a monthly basis. Profit rate is 14%. Required:

1.    What mode of finance will be used in this case? Write down the all steps and agreements which will be involved in this transaction.

2.    Provide the cost of 01 unit, rent per unit, monthly payments for first 05 months (rent + unit cost).

3.    Who will be the responsible for any damages happened with the house during the financing period and why? 


1
Expert's answer
2021-06-02T12:04:11-0400

1) Profit-and-loss sharing partnership (mudarabah)

Mudarabah is a profit-and-loss partnership arrangement in which one party (financier or rab-ul mal) distributes capital to another partner (labor provider or mudarib) who is responsible for capital management and investment. Profits are divided among the parties in accordance with a predetermined ratio.


2) "A=PRT +P"

"25-4=P \times0.14\times(5\times12)+P"

"21=8.4P +P"

"21=9.4P"

"P= Rs 2.23 M"


Cost per unit"= Rs2.23 *10^{-3} M"


3) Mr. Javed is labor provider or mudarib. He is responsible for any damages happened with the house during the financing period.